OpenStudy (anonymous):

what expression equals to ^3√24x^7y^6

OpenStudy (anonymous):

\[\sqrt[3]{24}x ^{7}y ^{6}\]

OpenStudy (anonymous):

is there a list of options?

OpenStudy (anonymous):

\[2x ^{3}y ^{3}\sqrt{3x}\]

OpenStudy (anonymous):

\[6xy ^{3}\sqrt{2xy}\]

OpenStudy (aum):

I suspect the cube root in the original problem covers all items and not just 24.

OpenStudy (anonymous):

\[x ^{6}y ^{63}\sqrt{36}\]

OpenStudy (aum):

Is it \(\Large \sqrt[3]{24}x ^{7}y ^{6} \) or \(\Large \sqrt[3]{24x ^{7}y ^6} \) ?

OpenStudy (anonymous):

those aren't on the answer choices :/

OpenStudy (aum):

No I am asking about the question.

OpenStudy (anonymous):

OH HAHA. yeah i got it lmao its the first one

OpenStudy (anonymous):

its \[\sqrt[3]{24x ^{7}}y ^{6}\]

OpenStudy (aum):

\( \Large \sqrt[3]{24x ^{7}} * y ^{6} = \sqrt[3]{8*3*x*x ^{6}} *y ^{6} = \sqrt[3]{(2x^2)^3*3x}*y ^{6} = \\ \Large 2x^2\sqrt[3]{3x}*y ^{6} \\ \Large \)

OpenStudy (aum):

Can you post a screenshot of the question?

OpenStudy (aum):

It was as I suspected all along. It was the second alternative I asked you about earlier. The question is: \(\Large \sqrt[3]{24x ^{7}y ^6}\) \( \Large \sqrt[3]{24x ^{7}y^6} = \sqrt[3]{8*3*x*x ^{6}y^6} = \sqrt[3]{(2x^2y^2)^3*3x} = \\ \Large 2x^2y^2\sqrt[3]{3x} \\ \Large \)

OpenStudy (driftracer305):

yea i agree....... just saw y^6...... cuberoot it is y^2..... simple easy 2 seconds...lol